Fermi National Laboratory
 

Accelerator Update for September 8, 2006 - September 11, 2006

 

Friday September 8

The day shift began with Operations monitoring store 4942 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with the Meson MTest experiment and MiniBooNE taking beam, and with NuMI off for target replacement.

Tevatron (TeV) experts began a Tevatron Electron Lens #2 (TEL2) end-of-store study at 1:08 PM.

At 3:03 PM, the TEL2 studies induced a quench at TeV sector E1. Investigations discovered that a vacuum burst occurred just before the quench. Experts investigated further.

Operations put the TeV into a dry squeeze at 5:38 PM. During the squeeze the TeV aborted due to an A0 coasting beam safety system valve permit. However, there was no valve problem and the abort reset immediately. Controls personnel replaced hardware cards in a CAMAC and vacuum crate.

Operations disabled the M-Test beam around 6:17 PM.

Saturday September 9

The midnight shift began with Operations monitoring the stash and stack, and with MiniBooNE taking beam.

Operations put the TeV into shot setup at 12:10 AM.

Operations established store 4947 at 2:26 AM with an initial luminosity of 206.18E30.

Store 4947 set a new luminosity record.

Operations established beam to M-Test at 6:39 AM.

At 12:30 PM, a Booster kicker (MKS07) tripped off. System experts came in and repaired the device. Beam operations returned at 3:20 PM.

Operations disabled the M-Test beam at 4:18 PM.

Booster kicker MKS07 failed again at 7:52 PM. The system experts came in and replaced the same device as before (PFN). Beam operations resumed by 9:33 PM.

At 10 PM, controls software problems became evident. Experts found that a name server initiated the problem.

Sunday September 10

The midnight shift began with Operations monitoring store 4947 and the stash, with Pbar stacking, with MiniBooNE taking beam, and with controls problems continuing. Numerous Java engines were found to have caused the controls issues.

Operations began sending beam to MTest at 7 AM.

Operations terminated the store 4947 at 10:05 AM. Shot setup was put on hold while controls personnel continued their work.

Operations put the TeV into shot setup at 2:30 PM.

Operations established store 4949 at 4:40 PM with an initial luminosity of 228.1E30.

Store 4949 set another luminosity record.

Operations disabled the MTest beam at 5:20 PM.

Monday September 11

The midnight shift began with Operations monitoring store 4949 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

Spill was established to M-Test at 6:56 AM.

At 8:54 AM, a Recycler corrector power supply tripped off, causing the loss of a large part of the stash.

The Plans for Monday

The plans for today are to keep store 4949 colliding in the TeV until this evening or tomorrow morning, depending on how soon the Recycler experts can find and fix the corrector problem.

NuMI will resume taking beam today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/12/2006    email Fermilab